OpenStudy (anonymous):

Find equations of the tangent lines to the curve y = (x − 1)/(x + 1) that are parallel to the line x − 2y = 3.

myininaya (myininaya):

find the slope of x-2y=3 find y' given y=(x-1)/(x+1) we want the slope of the linear equation to have the same slope as a tangent line of y=(x-1)/(x+1) so do the following set slope of linear line=y' and solve for x

OpenStudy (anonymous):

im confused the answer I need to find has 2 equations; each with a different y-intercept

myininaya (myininaya):

so what did you get for the slope of the linear equation above?

OpenStudy (anonymous):

1/2

myininaya (myininaya):

ok what did you get when finding y' given y=(x-1)/(x+1)?

OpenStudy (anonymous):

\[2/(x+1)^2\]

myininaya (myininaya):

[(x+1)-(x-1)]/(x+1)^2 yes i agree gj now set this =1/2 and solve for x

myininaya (myininaya):

2/(x+1)^2=1/2

myininaya (myininaya):

2/4=1/2 right?

myininaya (myininaya):

so we want (x+1)^2=4?

myininaya (myininaya):

so x+1=2 or x+1=-2

myininaya (myininaya):

=> we have the curve has the same slope as x-2y=3 at x=1 or x=-3

myininaya (myininaya):

i will find one equation for you you can do the other we have two equations one for x=1 and one for x=-3 ------------ IF x=1, then y=(x-1)/(x+1)=(1-1)/(1+1)=0 so we know a point on the tangent line (1,0) remember slope is 1/2 so we have the tangent equation at x=1 has form y=mx+b where m=1/2 and we can find b the y-intercept since we know a point (x=1,y=0) pluggin' this in we get 0=1*1/2+b -1/2=b so the equation at x=1 with slope 1/2 is y=x/2-1/2

myininaya (myininaya):

so you can do it for x=-3

OpenStudy (anonymous):

I'm getting more confused as you go on haha

myininaya (myininaya):

are you confused with the calculus part or the algebra part?

myininaya (myininaya):

it has to be the algebra part since we already done the cal part

myininaya (myininaya):

if we know a point on a line and the slope, then we can find the equation of the line right? we have the slope m=1/2 and we have the point (1,0) just plug into y=mx+b and solve for b then we have it

myininaya (myininaya):

well we have a point on the line*